Determine the complex number z which satisfies the equations |z + 3i| = |z + 5 − 2i| and |z − 4i| = |z + 2i| simultaneously.
1
Expert's answer
2020-05-20T17:39:25-0400

"\begin{cases}
 | z+3i| = | z+5-2i| \\
 | z - 4i | = | z + 2i |
\end{cases}"

"z=x+iy"

"|x+iy+3i| =|x+iy+5-2i|"

"\sqrt{x^2+(y+3)^2}=\sqrt{(x+5)^2+(y-2)^2}"

"x^2+y^2+6y+9=x^2+10x+25+y^2-4y+4"

"10y=10x+20\implies y=x+2"


"| z - 4i | = | z + 2i |"

 "|x+iy-4i| =|x+iy+2i|"

"\sqrt{x^2+(y-4)^2}=\sqrt{(x^2+(y+2)^2}"

"x^2+y^2-8y+16=x^2+y^2+4y+4"

"-8y+16=4y+4\implies12y=12\implies y=1"

"\begin{cases}
 y=x+2 \\
 y=1
\end{cases}"

"x=-1"    "y=1"

"z=i-1"